If the left hand side of the equation x2 - y 2 + x - 3 y + sec q = 0 can be factorized into two linear
factors then the value of q is
(a)pie/3
b) 7 pie/6
c) 4 pie / 3
d)5 pie / 6

Grade:11

1 Answers

vikas askiitian expert
13 years ago

x2 + x + -( y2 + 3y) + secq = 0

x2 + x +1/4 -1/4 -(y2+3y+9/4 - 9/4) + secq = 0                   

(x+1/2)2 - [(y+3/2)2 -9/4] + secq - 1/4 = 0

(x+1/2)2 - (y+3/2)2 + 9/4-1/4+secq = 0

this eq will factorize into two linear factors if the constant term is 0

secq + 2 = 0

cosq = -1/2

q = 2pi/3 , 4pi/3

option c is correct